# SRAGAN: Saliency Regularized and Attended Generative Adversarial Network for Chinese Ink-wash Painting Style Transfer

### v2

@@ 3.3.1 Saliency Attended Adversarial Loss
Taking the $\mathcal{X}\rightarrow\mathcal{Y}$ mapping as example, let $D^{1}_{\mathcal{Y}}(\cdot)$ denote the output $\frac{1}{16}$-scale logit matrix at the main branch of $D_{\mathcal{Y}}$, $D^{2}_{\mathcal{Y}}(\cdot)$ denote the predicted $\frac{1}{8}$-scale logit matrix at the auxiliary branch, $\mathbbm{1}^{1}$ be an all-one matrix of the same spatial size as $D^{1}_{\mathcal{Y}}(\cdot)$, $\mathbbm{1}^{2}$ be an all-one matrix of the same size as $D^{2}_{\mathcal{Y}}(\cdot)$, $S(\cdot)$ and $T(\cdot)$ denote the saliency detection network and the binary thresholding operation respectively, $\downarrow$ denotes ×$\times 8$ bilinear downsampling, then:
@@ 3.3.3 Saliency IOU Loss
in which $x^{\prime}$ is the painting generated by $G$ given the input real picture $x$, i.e., $x^{\prime}=G(x,(S(x)))$, $\hat{x}$ is the recovered result of $x$ inverted by the backward generator $F$, i.e., $\hat{x}=F(G(x,S(x)),S(x))$. Likewise, $y^{\prime}=F(y,(S(y)))$, $\hat{y}=G(F(y,S(y)),S(y))$. We resort to IOU-based saliency consistency constraint to enforce only overall object structure similarity, rather than strictly enforcing per-pixel consistency between saliency maps. Such relaxed structural constraint allows mild contour displacement in the stylized paintings and thus enables the network to focus on synthesizing more delicate and vivid artistic brush strokes and textures, producing style patterns that are more in line with the specific drawing skills of Chinese ink-wash painting. As visually verified in the experiment section, our IOU-based saliency consistency loss avoids rigid stylization effects caused by pixel-level saliency consistency constraint.
@@ 4.2 Training Details
For all the three Chinese ink-wash painting stylization tasks, the source-domain real pictures are partitioned into a training set and a test set in the proportion of 4:1. All images are resized to 256 $\times$ 256. we set $\lambda_{1}=1$, $\lambda_{2}=10$, $\lambda_{3}$ in Eq. 12, Two Adam optimizers with momentum parameters $\beta_{1}=0.5$, $\beta_{2}=0.999$ are applied to train the generators ($G$ and $F$) and the discriminators ($D_{\mathcal{X}}$ and $D_{\mathcal{Y}}$) respectively. We train for 200 epochs with the learning rate fixed at 0.0002 in the first 100 epochs and then linearly decayed to 0 in the next 100 epochs.
@@ 4.3 Qualitative Evaluation
Besides, we also compare to advanced diffusion model based I2I stylization methods including Null-text inversion [22], PnP [23], FreeStyle [24], and SDEdit [21]. These methods are based on pre-trained large-scale text-to-image diffusion model, for which we use Stable Diffusion v1.5 [20] as the foundation model. To stylize an input image into a Chinese ink-wash painting, we use “Chinese ink-wash painting of *” as text prompt where {``$a\ horse",``birds",``landscape"\}$. Example qualitative results displayed in Fig. 11 show that SDEdit with high denoising strength (e.g., SDEdit(0.75)) fails to maintain object content structure of the input image, while the counterpart with low denoising strength (e.g., SDEdit(0.4)) preserves content structure well at the expense of insufficient ink-wash style rendering. Results of Freestyle suffer from noticeable object structure distortion. In terms of stylization effect, all the compared diffusion-based methods produce results with less resemblance to the real style patterns of Chinese ink-wash paintings. By contrast, our results exhibit more precise Chinese ink-wash painting styles without sacrificing object content integrity.
@@ 4.4 Quantitative Evaluation
Table 1: Quantitative comparison to related NST and GAN-based methods.
@@
00footnotetext: Note: The “Landscape”, “Horse”, “Bird” denote the test set of the RealLandscape, RealHorses, RealBirds dataset, respectively.

### v3

@@ 3.3.1 Saliency Attended Adversarial Loss
Taking the $\mathcal{X}\rightarrow\mathcal{Y}$ mapping as example, let $D^{1}_{\mathcal{Y}}(\cdot)$ denote the output $\frac{1}{16}$-scale logit matrix at the main branch of $D_{\mathcal{Y}}$, $D^{2}_{\mathcal{Y}}(\cdot)$ denote the predicted $\frac{1}{8}$-scale logit matrix at the auxiliary branch, $\mathbbm{1}^{1}$ be an all-one matrix of the same spatial size as $D^{1}_{\mathcal{Y}}(\cdot)$, $\mathbbm{1}^{2}$ be an all-one matrix of the same size as $D^{2}_{\mathcal{Y}}(\cdot)$, $S(\cdot)$ and $T(\cdot)$ denote the saliency detection network and the binary thresholding operation respectively, $\downarrow$ denotes $\times 8$ bilinear downsampling, then:
@@ 3.3.3 Saliency IOU Loss
in which $x^{\prime}$ is the painting generated by $G$ given the input real picture $x$, i.e., $^{\prime}=G(x,(S(x))), x^\hat{x}$ is the recovered result of $x$ inverted by the backward generator $F$, i.e., $\hat{x}=F(G(x,S(x)),S(x))$. Likewise, $y^{\prime}=F(y,(S(y)))$, $\hat{y}=G(F(y,S(y)),S(y))$. We resort to IOU-based saliency consistency constraint to enforce only overall object structure similarity, rather than strictly enforcing per-pixel consistency between saliency maps. Such relaxed structural constraint allows mild contour displacement in the stylized paintings and thus enables the network to focus on synthesizing more delicate and vivid artistic brush strokes and textures, producing style patterns that are more in line with the specific drawing skills of Chinese ink-wash painting. As visually verified in the experiment section, our IOU-based saliency consistency loss avoids rigid stylization effects caused by pixel-level saliency consistency constraint.
@@ 4.2 Training Details
For all the three Chinese ink-wash painting stylization tasks, the source-domain real pictures are partitioned into a training set and a test set in the proportion of 4:1. All images are resized to 256 $\times$ 256. we set $\lambda_{1}=1$, $\lambda_{2}=10$, $\lambda_{3}=5$ in Eq. 12, Two Adam optimizers with momentum parameters $\beta_{1}=0.5$, $\beta_{2}=0.999$ are applied to train the generators ($G$ and $F$) and the discriminators ($D_{\mathcal{X}}$ and $D_{\mathcal{Y}}$) respectively. We train for 200 epochs with the learning rate fixed at 0.0002 in the first 100 epochs and then linearly decayed to 0 in the next 100 epochs.
@@ 4.3 Qualitative Evaluation
Besides, we also compare to advanced diffusion model based I2I stylization methods including Null-text inversion [22], PnP [23], FreeStyle [24], and SDEdit [21]. These methods are based on pre-trained large-scale text-to-image diffusion model, for which we use Stable Diffusion v1.5 [20] as the foundation model. To stylize an input image into a Chinese ink-wash painting, we use “Chinese ink-wash painting of *” as text prompt where $\in\{``a\ horse",``birds",``landscape$"\}. Example qualitative results displayed in Fig. 11 show that SDEdit with high denoising strength (e.g., SDEdit(0.75)) fails to maintain object content structure of the input image, while the counterpart with low denoising strength (e.g., SDEdit(0.4)) preserves content structure well at the expense of insufficient ink-wash style rendering. Results of Freestyle suffer from noticeable object structure distortion. In terms of stylization effect, all the compared diffusion-based methods produce results with less resemblance to the real style patterns of Chinese ink-wash paintings. By contrast, our results exhibit more precise Chinese ink-wash painting styles without sacrificing object content integrity.
@@ 4.4 Quantitative Evaluation
Table 1: Quantitative comparison to related NST and GAN-based methods.00footnotetext: Note: The “Landscape”, “Horse”, “Bird” denote the test set of the RealLandscape, RealHorses, RealBirds dataset, respectively.